OpenStudy (anonymous):

14x-(8x-9)=39 PLEASE HELP I KEEP GETTING THIS ANSWER WRONG

OpenStudy (anonymous):

ok lets go through it

OpenStudy (anonymous):

14x-8x+9=39

OpenStudy (anonymous):

13x-8x-30=0

OpenStudy (anonymous):

5x=30

OpenStudy (anonymous):

x=6

OpenStudy (anonymous):

did u get that?

OpenStudy (anonymous):

yes

OpenStudy (jdoe0001):

\(\bf 14x-(8x-9)=39\\ \quad \\ -(8x-9)\implies -1 \times 8x -1\times -9 \implies -8x+9\\ \quad \\ 14x-(8x-9)=39\implies 14x-8x+9=39\)